# Instruction-Guided Scene Text Recognition

### v2

@@ 3.2 IGTR Architecture and Learning
Image encoder. On the image side, SVTR-B backbone [5] (with the rectification module and CTC decoder removed) is utilized as the encoder. Given a text image of size $H\times W\times 3$, the visual features ($F_{v}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{4}\times D}$) are extracted by SVTR-B [5]. Subsequently, image embeddings ($E_{I}\in\mathbb{R}^{\frac{HW}{32}\times D}$) are obtained by flattening the height ($H$) and width ($W$) of $F_{v}$. Here, $D=384$ represents the dimension of the features.
@@
Instruction encoder. The instruction encoder aims to extract distinct features for each instruction. Since both condition and question-answer are composed of character attributes, we can use character attributes to describe the instruction. Therefore, unlike previous methods in the natural image domain employ pre-trained text encoders [48, 20] to encode the instruction, which is computationally costly, we devise a lightweight encoder to map the instruction to two embeddings, i.e., condition, and question embeddings.
@@
We first inspect the compositional elements of the character attributes, which consist of five basic attribute elements, i.e., character, frequency, position, status, and constraint. All are enumerable. Therefore, as depicted in Fig. 3, the instruction encoder consists of five learnable embedding layers. They are character embedding ($CE\in\mathbb{R}^{C\times D}$), frequency embedding ($FE\in\mathbb{R}^{F\times D}$), position embedding ($PE\in\mathbb{R}^{P\times D}$), status embedding ($SE\in\mathbb{R}^{2\times D}$), and constraint embedding ($ConsE\in\mathbb{R}^{L\times D}$) layers. In these embeddings, $C$, $P$, and $F$ signify the size of the character set, the maximum position index, and the maximum character frequency, respectively. we set $P$, $F$ and $L$ to the same length. For each attribute element, the encoding begins by mapping it to the index in the corresponding set. Then, its embedding is found by looking up the vector at this index in the corresponding embedding layer.
@@
We then calculate the representation of each condition and question using attribute elements. For those associated with only one attribute element, they are directly represented by the embedding of that element. For instance, “What is the 4-th character?” is described by a 384-d embedding from $PE$. Note that some condition and question are related to multiple attribute elements, e.g., ce, cf and pc. Their representations are obtained by summing the embeddings of its constituent elements. As illustrated in Fig. 3, “The 1st character is A” can be written as 1-A, and the embedding is the sum of the corresponding vectors from $CE$ and $PE$. Moreover, for questions related to $\texttt{cf}_{cons}$, the embedding is further added with a constraint embedding from $ConsE$. In addition, since character position and frequency questions (see Tab. I) are both described by the unitary attribute embedding, a position token ($pt\in\mathbb{R}^{1\times D}$) and a frequency token ($st\in\mathbb{R}^{1\times D}$) are also defined. They are added to the corresponding unitary attribute embedding for differentiation. As the example, “How many times does A appears?” can be described by an embedding that sums $st$ and the corresponding vector from $CE$, and “Which character in the image is R?” is characterized by an embedding summed $pt$ from the vector in $CE$.
@@
With the embeddings above, we generate the partition-level condition and question embeddings. The former, i.e., $E_{c}\in\mathbb{R}^{L_{c}\times D}$, is obtained by concatenating all condition embeddings in that partition. While the latter is obtained by grouping the question embeddings into four sets according to answer types, i.e., $E_{q}$ = [$E_{qc}$;$E_{qf}$;$E_{qp}$;$E_{qs}$] $\in\mathbb{R}^{L_{q}\times D}$, where $E_{qc}$ corresponds to question set $Q_{c}$ and the rest similarly defined. $L_{q}$ and $L_{c}$ are the number of Ques and Cond. Although a large number of instructions are sampled given a text instance, the K partitions and the questions in each partition are independent to each other. They thus can be processed in parallel, which enjoys efficient computation with almost no increase in training cast.
@@
Cross-modal feature fusion. Once instructions and images are successfully encoded, effectively fusing them is crucial for multi-modal tasks. A common practice involves treating the instruction embedding as the Query and the image embedding as the Key and Value, performing cross-attention to extract relevant information from the image modality. Since the instruction in Tab. I includes condition that represents the known fact, and question that denotes the asked question. Thus, it is necessary to construct a new module to enable the effective interaction and fusion of these three embeddings.
We devise a dedicated cross-modal feature fusion (CMFF) module to fuse question ($E_{q}$), condition ($E_{c}$) and image embeddings ($E_{I}$). First, with $E_{c}$ as Key and Value, and each of the three embedding as Query, three multi-head attentions (MHA, Eq. 1) [31] are employed. This process simultaneously performs self-attention on condition and cross-attention involving $E_{c}$ and the other two kinds of embeddings, fully integrating $E_{c}$ with both other modalities and itself. Second, the same process is performed again with $E_{I}$ switch $E_{c}$ as Key and Value. Consequently, the three modalities further absorb desired information from the image side, effectively enabling IGTR to understand the known attributes and the image content.
@@
In Eq. 1, there are $h$ attention heads. For the i-th head, it associates with three different weight matrices $W_{i}^{q}\in\mathbb{R}^{D\times D_{h}},W_{i}^{k}\in\mathbb{R}^{D\times D_{h}},W_{i}^{v}\in\mathbb{R}^{D\times D_{h}}$, [;] denotes the concatenation, $D_{h}$ = $\frac{D}{h}$, and $W^{o}\in\mathbb{R}^{D\times D}$.
Notably, as shown in the upper right portion of Fig. 2, three MHA modules are involved in each interaction, and their parameters are shared to ensure compact and fast inference of IGTR. In this way, two interactions of the three modalities can be described by Eq. 2:
@@
In the following process (Eq. 3), the $E_{q}^{2}$ from Eq. 2 is employed as Query, and question-to-condition and question-to-image cross-attentions are successively conducted, from which the question extracts finer features from both condition and the image domain to obtain answer-related cross-modal embeddings $\hat{E}_{q}$.
$E_{q}^{j_{1}}$ $=LN(\text{MHA}(E_{q}^{i},E_{c}^{2},E_{c}^{2})+E_{q}^{j})$
$E_{q}^{j_{2}}$ $=LN(FFN(E_{q}^{j_{1}})+E_{q}^{j_{1}})$
$E_{q}^{j_{3}}$ $=LN(\text{MHA}(E_{q}^{j_{2}},E_{I}^{2},E_{I}^{2})+E_{q}^{j_{2}})$
$E_{q}^{j+1}$ $=LN(FFN(E_{q}^{j_{3}})+E_{q}^{j_{3}})$ (3)
where $j$ starts at 2, $E_{q}^{j_{*}}$ are the intermediate variables, and $E_{q}^{j+1}$ denoted as $\hat{E}_{q}$ is the output of the CMFF.
Multi-task answer head. $\hat{E}_{q}$ represents the question embedding to be decoded and it can be split into $\hat{E}_{qc}\in\mathbb{R}^{N_{c}\times D}$, $\hat{E}_{qf}\in\mathbb{R}^{N_{f}\times D}$, $\hat{E}_{qp}\in\mathbb{R}^{N_{p}\times D}$, and $\hat{E}_{qs}\in\mathbb{R}^{N_{s}\times D}$ based on the four answer types. With those embeddings, a multi-task answer head is constructed to perform answer prediction tasks. Specifically, the character head, frequency head, status head, and position head are responsible for instructions whose answer is the character, frequency, status (Yes or No), and position index, respectively. The four heads are all linear classifiers, each associated with learnable parameter matrices: $W_{c}\in\mathbb{R}^{D\times C}$, $W_{f}\in\mathbb{R}^{D\times F}$, $W_{p}\in\mathbb{R}^{D\times 1}$, and $W_{s}\in\mathbb{R}^{D\times P},$ respectively. Consequently, the prediction answers $\hat{A}_{c}$, $\hat{A}_{f}$, $\hat{A}_{p}$, and $\hat{A}_{s}$ are obtained applying the four heads to $\hat{E}_{q}$:
@@
where $N_{c}$, $N_{f}$, $N_{p}$, and $N_{s}$ are the number of answers corresponding to the four types, $L_{q}=N_{c}+N_{f}+N_{p}+N_{s}$, and $\sigma_{1}$ and $\sigma_{2}$ are softmax and sigmoid functions.
Answer loss. Prediction results of the four heads are compared with the answer label to calculate loss. Assuming $\hat{A}^{n}_{c}$ and ${A}^{n}_{c}$ are the prediction and answer label of the n-th question in $Q_{c}$ and the rest defined similarity, the four losses are calculated as follows and then summed to get a total loss $\mathcal{L}$ according to the Eq. 4:
@@
Figure 5: Illustration of using the two questions described in Sec. 4.2 for zero-shot text inference of image MOMS. The $condition$ is set to None.
@@ 3.3 Text Recognition with instruction
When the model is sufficiently trained with character attribute prediction instructions, it is deemed to understand character attributes and the text image. Then, recognition instructions as shown in Tab. II are employed for training, where each row corresponds to a different recognition pipeline. Specifically, the first two rows delineate the well-known PR [2, 3] and AR [8, 10] pipelines, standardized in an instruction-guided manner. Similar to existing methods [2, 3, 8, 10], both PR and AR are limited to processing text whose length is up to 25 characters. The third row introduces re-identification (RI). This instruction can guide the model to double-check the i-th character, which is particularly useful when it is combined with PR or AR to recognize previously misidentified characters. In the fourth row we propose extrapolating recognition (ER), a novel pipeline analogous to AR. Starting from a given sub-string, each time it infers the previous and next characters of the sub-string, thus enabling a sub-string-based stepwisely extending recognition. It offers an elegant way to recognize text whose length is beyond the previous limit of 25 characters. In Fig. 4, we illustrate the recognition procedure by using the four types of instructions, aiding in elucidating their operational principles and implementation nuances.
In addition, benefiting from the flexibility of the instruction-guided paradigm, text recognition can also be achieved by training on character attribute prediction instructions alone (as detailed in Fig. 5), devoid of recognition instructions. Similar to [20, 12, 13, 14], this pipeline is denoted as zero-shot recognition, whose performance we will verify in the experiments.
TABLE III: Ablations on instruction variants.
@@ 4.1 Datasets and Implementation Details
We evaluate IGTR on both English and Chinese datasets. For English our models are trained on MJSynth (MJ) [49, 50] and SynthText (ST) [51], two widely used synthetic scene text datasets. Then the models are tested on: (1) six regular and irregular text benchmarks, i.e., ICDAR 2013 (IC13) [52], Street View Text (SVT) [53], IIIT5K-Words (IIIT) [54], ICDAR 2015 (IC15) [55], Street View Text-Perspective (SVTP) [56] and CUTE80 (CUTE) [57]. For IC13 and IC15, we use the versions with 857 and 1,811 images, respectively. We abbreviate the six benchmarks as common benchmarks. (2) Real-world Union14M-L benchmark. It contains over 3.2 million training images and over 0.4 million test images. All are real-world images with both complexity and versatility [58]. The test set includes seven challenging subsets including curve, multi-oriented, artistic, etc. For Chinese, we use Chinese text recognition (CTR) dataset [59], a benchmark containing four subsets: Scene, Web, Document, and Writing. We train the model on the whole training set and use the validation subset of Scene to determine the best model, which is then assessed on the four test subsets.
@@
We use AdamW optimizer [60] with a weight decay of 0.05 for training. For English models, all images are resized while maintaining their aspect ratio, with a maximum pixel count of $32\times 128$ [61, 11, 34]. The learning rate (LR) is set to $5\times 10^{-4}$ and batchsize is set to 768. One cycle LR scheduler [62] with 1.5 epochs linear warm-up is used in all the 20 epochs. The same as [61, 5, 58], data augmentation like rotation, perspective distortion, motion blur and gaussian noise, are randomly performed during training. The alphabet includes all case-insensitive alphanumerics. For Chinese models, all text instances are resized to $32\times 256$ and data augmentation is not performed following [63]. The LR is set to $5\times 10^{-4}$ and batchsize is set to 512. One cycle LR scheduler with 3 epochs linear warm-up is used in all 100 epochs. Word accuracy is used as the evaluation metric. The size of the character set $C$ is set to 37 for English and 6625 for Chinese [23]. The maximum prediction length $L$ is set to 25 for both. All models are trained with mixed-precision on 2 Tesla A100 GPUs.
@@
TABLE IV: Ablations on CMFF components, where recognition accuracy (%) and inference time (ms) are both given.
Model PR AR ER
Acc Time Acc Time Acc Time
Full Model 82.51 3.98 84.86 10.3 83.78 9.52
(a): w/o Query to Condition 81.75 3.91 82.41 9.87 82.13 9.36
(b): w/o Query to Image 81.85 3.87 82.35 9.54 82.02 9.21
(c): (a) + (b) 80.85 3.77 81.23 9.35 80.83 9.06
(d): w/o Question to Condition 81.96 3.93 82.22 9.86 82.01 9.39
(e): w/o Question to Image 81.34 3.90 83.21 9.53 82.88 9.24
(f): (d) + (e) 80.93 3.81 82.01 9.37 81.25 9.12

### v3

@@ 3.2 IGTR Architecture and Learning
Image encoder. On the image side, SVTR-B backbone [5] (with the rectification module and CTC decoder removed) is utilized as the encoder. Given a text image of size $H\times W\times 3$, visual features ($F_{v}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{4}\times D}$) are extracted by SVTR-B [5]. Subsequently, the image embedding ($E_{I}\in\mathbb{R}^{\frac{HW}{32}\times D}$) is obtained by flattening the height ($H$) and width ($W$) of $F_{v}$. Here, $D=384$ represents the dimension of the features.
@@
Instruction encoder. The instruction encoder aims to extract distinct features for each instruction. Since both condition and question-answer are composed of character attributes, we can describe the instruction using attribute-level representation. Therefore, unlike previous methods that employ pre-trained text encoders [47, 20] to compile the instruction, which is computationally costly, we devise a lightweight encoder to map the instruction to condition and question embeddings.
@@
We first inspect the compositional elements of the character attributes, which consist of five basic attribute elements, i.e., character, frequency, position, status, and constraint. All are enumerable. Therefore, as depicted in Fig. 3, the instruction encoder consists of five learnable embedding layers. They are character embedding ($CE\in\mathbb{R}^{C\times D}$), frequency embedding ($FE\in\mathbb{R}^{F\times D}$), position embedding ($PE\in\mathbb{R}^{P\times D}$), status embedding ($SE\in\mathbb{R}^{2\times D}$), and constraint embedding ($ConsE\in\mathbb{R}^{L\times D}$) layers. In these embeddings, $C$, $F$, and $P$ signify the size of the character set, the maximum character frequency, and the maximum position index, respectively. We set $F$, $P$ and $L$ to the same value. For each attribute element, the encoding begins by mapping it to the index in the corresponding set. Then, its embedding is found by looking up the vector at this index in the corresponding embedding layer.
@@
We then represent each condition and question using attribute elements. For those associated with only one attribute element, they are directly represented by the embedding of that element. For instance, “What is the 4-th character?” is described by a 384-d embedding denoting 4 from $PE$. Note that condition and some question are related to multiple attribute elements, e.g., ce, cf and pc. Their representations are obtained by summing the embeddings of its constituent elements. As illustrated in Fig. 3, “The 1st character is A” can be written as 1-A, and the embedding is the sum of the corresponding vectors from $PE$ and $CE$. Moreover, for questions related to $\texttt{cf}_{cons}$, the embedding is further added with a constraint embedding from $ConsE$. To precisely describe a substring, we first introduce an order token ($ot\in\mathbb{R}^{1\times D}$) to distinguish the order of characters within a substring. Then, each character in the substring is represented by the sum of its character embedding and corresponding order token, and the substring is represented by further summing the obtained embeddings. In addition, since character position and frequency questions (see Tab. I) are both described by unitary attribute embedding, a position token ($pt\in\mathbb{R}^{1\times D}$) and a frequency token ($ft\in\mathbb{R}^{1\times D}$) are also defined. They are added to the corresponding unitary attribute embedding for differentiation. As the example, “How many times does A appears?” can be described by an embedding that sums $ft$ and the corresponding “A” vector from $CE$, and “Which character in the image is R?” is characterized by an embedding summed $pt$ from the vector denoting “R” in $CE$.
@@
With the mapping above, we get the partition-level condition and question embeddings. The former, i.e., $E_{c}\in\mathbb{R}^{L_{c}\times D}$, is obtained by concatenating all $L_{c}$ condition embeddings in that partition. While the latter is obtained by grouping the question embeddings into four sets according to answer types, i.e., $E_{q}$ = [$E_{qc}$;$E_{qf}$;$E_{qp}$;$E_{qs}$] $\in\mathbb{R}^{L_{q}\times D}$, where $E_{qc}$ corresponds to question set $Q_{c}$ and the rest similarly defined, [;] denotes feature concatenation, $L_{q}$ is the number of questions. Although a large number of instructions are sampled given a text instance, the K partitions and the questions in each partition are independent to each other. They thus can be processed in parallel, which enjoys efficient computation with almost no increase in training cast.
To enable parallel processing, we process the K partitions sequentially, predicting the four types of answers independently for each partition. This approach ensures consistent shapes for both condition and question embeddings across all partitions. Specifically, when the length of $E_{c}$, $E_{qc}$, or $E_{qs}$ falls short of $P$, they are padded with placeholder tokens to maintain consistent shapes. For $E_{qf}$ and $E_{qp}$, their shapes are both set to $C\times D$. The padded tokens are subsequently masked during the multi-head attention computation and ignored in loss calculations. As a result, this padding strategy facilitates seamless handling of variable-shape embeddings as inputs and ensures efficient parallel training within Transformer-based architectures.
@@
Cross-modal feature fusion. Once the image and instructions are encoded, effectively fusing them is crucial for multi-modal tasks. A common practice involves treating the instruction embedding as the Query and the image embedding as the Key and Value, performing cross-attention to extract relevant information from the image modality. Since the instruction in Tab. I includes condition that represents the known fact, and question that denotes the asked question, it is necessary to construct a new module to enable the effective interaction and fusion of these three embeddings.
We devise a dedicated cross-modal feature fusion (CMFF) module to fuse question ($E_{q}$), condition ($E_{c}$) and image embeddings ($E_{I}$). It consists of four cross-attention stages. The first is Query to Condition attentions. With $E_{c}$ as Key and Value, and all the three embedding as Query, three multi-head attentions (MHA, Eq. 1) [48] are employed. This stage simultaneously performs self-attention on $E_{c}$ and cross-attention on $E_{q}$ and $E_{c}$, and on $E_{I}$ and $E_{c}$. These interactions fully integrate $E_{c}$ with both other modalities and itself. The second is Query to Image attentions, the same process as above is performed again with $E_{I}$ switch $E_{c}$ as Key and Value. Consequently, the three modalities further absorb desired information from the image side, effectively enabling IGTR to understand the known attributes and image content.
@@
In Eq. 1, the head number $h$ is set to 12, $W^{o}\in\mathbb{R}^{D\times D}$. The each head associates with three different weight matrices $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{D\times D_{h}}$, where $D_{h}$ = $\frac{D}{h}$.
Notably, as shown in the upper right of Fig. 2, three MHA modules are involved in each interaction, and their parameters are shared to ensure compact architecture and fast inference of IGTR. In this way, the two interactions above can be described by Eq. 2:
@@
In the following process (Eq. 3), $E_{q}^{2}$ from Eq. 2 is employed as Query, and question-to-condition and question-to-image cross-attentions are successively conducted. They correspond to the third and fourth cross-modality interactions, from which the question extracts finer features from both condition and the image domain to obtain answer-related cross-modal embeddings $\hat{E}_{q}$.
$E_{q}^{3}$ $=LN(\text{MHA}(E_{q}^{2},E_{c}^{2},E_{c}^{2})+E_{q}^{2})$
$E_{q}^{3}$ $=LN(FFN(E_{q}^{3})+E_{q}^{3})$
$E_{q}^{4}$ $=LN(\text{MHA}(E_{q}^{3},E_{I}^{2},E_{I}^{2})+E_{q}^{3})$
$\hat{E}_{q}$ $=LN(FFN(E_{q}^{4})+E_{q}^{4})$ (3)
Multi-task answer head. $\hat{E}_{q}$ represents the question embedding to be decoded and it can be split into $\hat{E}_{qc}\in\mathbb{R}^{N_{c}\times D}$, $\hat{E}_{qf}\in\mathbb{R}^{N_{f}\times D}$, $\hat{E}_{qp}\in\mathbb{R}^{N_{p}\times D}$, and $\hat{E}_{qs}\in\mathbb{R}^{N_{s}\times D}$ based on the four answer types. $N_{c}$, $N_{f}$, $N_{p}$, and $N_{s}$ are the number of answers corresponding to the four types and $L_{q}=N_{c}+N_{f}+N_{p}+N_{s}$. With those embeddings, a multi-task answer head is constructed to perform answer prediction. Specifically, the character head, frequency head, position head, and status head are responsible for questions whose answer is the character, frequency, position index, and status (Yes or No), respectively. The four heads are all linear classifiers, each associated with learnable parameter matrices: $W_{c}\in\mathbb{R}^{D\times C}$, $W_{f}\in\mathbb{R}^{D\times F}$, $W_{p}\in\mathbb{R}^{D\times 1}$, and $W_{s}\in\mathbb{R}^{D\times P},$ respectively. Consequently, the prediction answers $\hat{A}_{c}$, $\hat{A}_{f}$, $\hat{A}_{p}$, and $\hat{A}_{s}$ are obtained by applying the four heads to $\hat{E}_{q}$:
@@
where $\sigma_{1}$ and $\sigma_{2}$ are softmax and sigmoid functions.
Answer loss. Prediction results of the four heads are compared with answer labels to calculate loss. Assuming $\hat{A}^{n}_{c}$ and ${A}^{n}_{c}$ are the prediction and answer label of the n-th question in $Q_{c}$ and the rest defined similarity, the four losses are calculated as follows and then summed to get a total loss $\mathcal{L}$ according to the Eq. 4:
@@ 3.3 Text Recognition with instruction
During model training, both the attribute prediction instructions in Tab. I and the recognition instructions in Tab. II are employed. The attribute prediction instructions guide IGTR in learning fine-grained character-level and relational features, fostering a nuanced understanding of textual patterns. On the other hand, the recognition instructions, each corresponding to a different recognition pipeline, focus on equipping IGTR with the specific mechanisms for recognizing text. They together can enable the enhanced text image understanding capability.
When the model is sufficiently trained, during inference IGTR utilizes either the parallel recognition (PR), autoregressive recognition (AR), re-identification (RI), or extrapolated recognition (ER) instructions in Tab. II to recognition text. Specifically, the first two instructions delineate the well-known PR and AR pipelines, respectively, standardized in an instruction-guided manner. Similar to existing methods [2, 3, 8, 10], both the two pipelines are limited to processing text whose length is up to 25 characters. The third instruction introduces RI. This instruction can guide the model to double-check the i-th character, which is particularly useful when it is combined with PR or AR to recognize previously misidentified characters. In the fourth row, we propose ER, a novel pipeline somewhat analogous to AR. Starting from a given sub-string, each time it infers the previous and next characters of the sub-string, thus enabling a sub-string-based stepwisely extending recognition. It offers an elegant way to recognize text whose length is beyond the previous limit of 25 characters. In Fig. 4, we illustrate the recognition procedures by using the four types of instructions, which aid in elucidating their respective operational principles and implementation details.
TABLE III: Ablations on instruction variants, where accuracy (%) of different recognition pipelines is given. The same below.
@@ 4.1 Datasets and Implementation Details
We evaluate IGTR on both English and Chinese datasets. For English the employed training datasets include: (1) MJSynth (MJ) [49, 50] and SynthText (ST) [51], the two widely used synthetic scene text datasets. (2) Real-world Union14M-L training set, which contains over 3.2 million labeled real-world text images with both complexity and versatility [52]. Note that Union14M-L training set overlaps with Union14M-L benchmark (test set) in 6,600 samples. These samples are filtered out from the training set to avoid data leakage. As for test protocols, the trained models are tested on: (1) six regular and irregular text benchmarks, i.e., ICDAR 2013 (IC13) [53], Street View Text (SVT) [54], IIIT5K-Words (IIIT) [55], ICDAR 2015 (IC15) [56], Street View Text-Perspective (SVTP) [57] and CUTE80 (CUTE) [58]. For IC13 and IC15, we use the versions with 857 and 1,811 images, respectively. We call the six benchmarks as Common benchmarks (abbreviated as Common). (2) Union14M-L benchmark [52] (abbreviated as Union14M-L). It contains over 0.4 million test images. The benchmark is composed of seven challenging subsets including curve, multi-oriented, artistic, etc. For Chinese, we use Chinese text recognition (CTR) dataset [59], a benchmark containing four subsets: Scene, Web, Document, and Writing. We train the model on the whole training set and use the validation subset of Scene to determine the best model, which is then assessed on the four test subsets.
@@
We use AdamW optimizer [60] with a weight decay of 0.05 for training. For English models, all images are resized while maintaining their aspect ratio, with a maximum pixel count of $32\times 128$ [61, 11, 33]. The learning rate (LR) is set to $5\times 10^{-4}$ and batchsize is set to 768. One cycle LR scheduler [62] with 1.5 epochs linear warm-up is used in all the 20 epochs. The same as [61, 5, 52], data augmentation like rotation, perspective distortion, motion blur and gaussian noise, are randomly performed during training. The alphabet includes all case-insensitive alphanumerics. For Chinese models, all text instances are resized to $32\times 256$ and data augmentation is not performed following [63]. The LR is also set to $5\times 10^{-4}$ and batchsize is set to 512. One cycle LR scheduler with 3 epochs linear warm-up is used in all 100 epochs. Word accuracy is used as the evaluation metric. The size of the character set $C$ is set to 37 for English and 6625 for Chinese [23]. The maximum prediction length $L$ is set to 25 for both. All models are trained with mixed-precision on 2 Tesla A100 GPUs.
@@
TABLE IV: Ablations on CMFF components, where recognition accuracy (%) and inference time (ms) are both given
. Model PR AR ER Acc Time Acc Time Acc Time Full Model 82.51 3.98 84.86 10.3 83.78 9.52 (a): w/o Query to Condition 81.75 3.91 82.41 9.87 82.13 9.36 (b): w/o Query to Image 81.85 3.87 82.35 9.54 82.02 9.21 (c): (a) + (b) 80.85 3.77 81.23 9.35 80.83 9.06 (d): w/o Question to Condition 81.96 3.93 82.22 9.86 82.01 9.39 (e): w/o Question to Image 81.34 3.90 83.21 9.53 82.88 9.24 (f): (d) + (e) 80.93 3.81 82.01 9.37 81.25 9.12
Figure 5: Two illustrations of using different attribute prediction instructions described in Sec. 4.3 for text inference of image MOMS. The $condition$ is set to None.